# Existence and non-existence of consistent estimators in supercritical controlled branching processes

### v1

@@ 5 Proofs
$||\mathcal{L}_{(Z_{j+1},\dots,Z_{k})|Z_{j}=u_{j}}-\mathcal{L}_{(X_{j+1},\dots,X_{k})|X_{j}=u_{j}}||_{TV}\leq\boldsymbol{\mathfrak{K}}_{k-j}(u_{j}),$
@@
$||\mathcal{L}_{(Z_{j},\dots,Z_{k})|Z_{j-1}=u_{j-1}}-\mathcal{L}_{(X_{j+1},\dots,X_{k})|X_{j-1}=u_{j-1}}||_{TV}$
$\leq||\mathcal{L}_{Z_{j}|Z_{j-1}=u_{j-1}}-\mathcal{L}_{X_{j}|X_{j-1}=u_{j-1}}||_{TV}+\mathbbm{P}(Z_{j}\leq N|Z_{j-1}=u_{j-1})+\boldsymbol{\mathfrak{K}}_{k-j}(N+1)$
$\leq||\mathcal{L}_{Z_{j}|Z_{j-1}=u_{j-1}}-\mathcal{L}_{X_{j}|X_{j-1}=u_{j-1}}||_{TV}+\frac{\varepsilon(u_{j-1})\cdot\sigma^{2}+\nu^{2}(u_{j-1})\cdot m^{2}}{(\varepsilon(u_{j-1})\cdot m-N)^{2}}+\boldsymbol{\mathfrak{K}}_{k-j}(N+1)$
for $u_{j-1}\geq M_{3}$ and $M_{3}\leq N<\varepsilon(u_{j-1})\cdot m$. Given (6) and that $||\mathcal{L}_{Z_{1}|Z_{0}=z}-\mathcal{L}_{X_{1}|X_{0}=z}||_{TV}=O(z^{-q})$, and taking $N:=\lfloor\alpha t\cdot u_{j-1}\rfloor$ for $\alpha\in(1/t,1)$, we can simplify the above bound to
$||\mathcal{L}_{(Z_{j},\dots,Z_{k})|Z_{j-1}=u_{j-1}}-\mathcal{L}_{(X_{j},\dots,X_{k})|X_{j-1}=u_{j-1}}||_{TV}$
@@
$\leq\frac{s}{u_{j-1}^{q}}+\frac{(a\sigma^{2}+bm^{2})}{(1-\alpha)^{2}t^{2}\cdot u_{j-1}}+\boldsymbol{\mathfrak{K}}_{k-j}(\alpha t\cdot u_{j-1})$
$=:\boldsymbol{\mathfrak{K}}_{k-j+1}(u_{j-1})$
Since we assumed that $\boldsymbol{\mathfrak{K}}_{k-j}$ was a decreasing function on $[M_{3},\infty)$, we see that $\boldsymbol{\mathfrak{K}}_{k-j+1}$ is also a decreasing function on $[M_{3},\infty)$.
Having shown the recursive relationship (16), and since $\alpha\in(1/t,1)$ implies $\alpha t>1$, by taking $f_{i}=\boldsymbol{\mathfrak{K}}_{i}$ for $i\in[k]$ we see that the sequence of functions $\{\boldsymbol{\mathfrak{K}}_{i}\}_{i\in[k]}$ satisfies the requirements of Lemma 5.2. It hence follows that
$\lim_{z\to\infty}\lim_{k\to\infty}\boldsymbol{\mathfrak{K}}_{k}(z)=0,$
but since
$0\leq||\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}=\lim_{k\to\infty}||\mathcal{L}_{(Z_{1},\dots,Z_{k})|Z_{0}=z_{0}}-\mathcal{L}_{(X_{1},\dots,X_{k})|X_{0}=z_{0}}||_{TV}\leq\lim_{k\to\infty}\boldsymbol{\mathfrak{K}}_{k}(z_{0}),$
it further follows that $\lim_{z_{0}\to\infty}||\mathcal{L}_{\{X_{n},z_{0}\}}-\mathcal{L}_{\{Z_{n},z_{0}\}}||_{TV}=0$. $\blacksquare$
Proof of Proposition 3.2. For a given set $\Pi$ of supercritical CBPs satisfying (6) and with transition probabilities parameterised by $\boldsymbol{\theta}$, assume there exist processes $\{Z_{n},z_{0}\},\{X_{n},z_{0}\}\in\Pi$ with $\boldsymbol{\theta}_{Z}\neq\boldsymbol{\theta}_{X}$ such that $||\mathcal{L}_{Z_{1}|Z_{0}=z_{0}}-\mathcal{L}_{X_{1}|X_{0}=z_{0}}||_{TV}=O(z_{0}^{-q})$ for some $q>0$. Let $\{\hat{\boldsymbol{\theta}}_{k}\}_{k\in\mathbbm{N}_{1}}$ be a sequence of estimators for $\boldsymbol{\theta}$.
Suppose that the sequence of estimators $\{\hat{\boldsymbol{\theta}}_{k}\}_{k\in\mathbbm{N}_{1}}$ is (weakly) consistent for $\boldsymbol{\theta}$ on the set of unbounded growth of the process for every initial population size. Then there exists a subsequence $\{k_{j}\}_{j\in\mathbbm{N}_{1}}$ such that $\{\hat{\boldsymbol{\theta}}_{k_{j}}\}_{j\in\mathbbm{N}_{1}}$ forms a strongly consistent sequence of estimators—that is, on the set of unbounded growth, $\hat{\boldsymbol{\theta}}:=\lim_{j\to\infty}\hat{\boldsymbol{\theta}}_{k_{j}}$ exists and equals $\boldsymbol{\theta}$ almost surely (see, for example, [6, Theorem 2.3.2]). In a slight abuse of notation, we say that $\hat{\boldsymbol{\theta}}$ is itself strongly consistent.
@@
{$||\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}=\sup_{C\subseteq\mathbbm{N}_{0}^{\infty}}|\mathbbm{P}(\{Z_{n},z_{0}\}\in C)-\mathbbm{P}(\{X_{n},z_{0}\}\in C)|,$
@@
{$||\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}\geq|\mathbbm{P}(\hat{\boldsymbol{\theta}}(\{Z_{n},z_{0}\})=\boldsymbol{\theta}_{X},\;Z_{n}\to\infty)-\mathbbm{P}(\hat{\boldsymbol{\theta}}(\{X_{n},z_{0}\})=\boldsymbol{\theta}_{X},\;X_{n}\to\infty)|.$ (17)
But, since $\hat{\boldsymbol{\theta}}$ is a strongly consistent estimator for $\boldsymbol{\theta}$ and $\boldsymbol{\theta}_{Z}\neq\boldsymbol{\theta}_{X}$,
$\mathbbm{P}(\hat{\boldsymbol{\theta}}(\{Z_{n},z_{0}\})=\boldsymbol{\theta}_{X},\;Z_{n}\to\infty)=0,$
while
$\mathbbm{P}(\hat{\boldsymbol{\theta}}(\{X_{n},z_{0}\})=\boldsymbol{\theta}_{X},\;X_{n}\to\infty)=\mathbbm{P}(X_{n}\to\infty).$
@@
{$||\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}\geq\mathbbm{P}(X_{n}\to\infty).$
@@
    Since $||\mathcal{L}_{Z_{1}|Z_{0}=z_{0}}-\mathcal{L}_{X_{1}|X_{0}=z_{0}}||_{TV}=O(z_{0}^{-q})$, by Lemma 3.1, $\lim_{z_{0}\to\infty}||\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}=0$,
@@
    Given (6), [8, Theorem 3.2] tells us that $\lim_{z_{0}\to\infty}\mathbbm{P}(Z_{n}\to\infty)=1$.
This creates a contradiction, since (i) and (ii) tell us that we can find $z_{0}\in\mathbbm{N}_{1}$ sufficiently large such that
{$||\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}<\mathbbm{P}(X_{n}\to\infty).$
$\blacksquare$
@@
Proof of Lemma 3.3. Since our two processes $\{Z_{n},z_{0}\}$ and $\{X_{n},z_{0}\}$ belong to the set $\Pi^{(k)}$, they must have a common control function, $\phi(\cdot)$. If we consider the realisations of $\phi(\cdot)$ as forming interstitial states within the processes $\{Z_{n},z_{0}\}$ and $\{X_{n},z_{0}\}$, we can recognise the expanded processes $\{\phi(z_{0}),\,Z_{1},\,\phi(Z_{1}),\,Z_{2},\,\dots\}$ and $\{\phi(z_{0}),\,X_{1},\,\phi(X_{1}),\,X_{2},\,\dots\}$, consisting of alternating progenitor numbers and population sizes, as two time-inhomogeneous Markov chains. We can see from the definition of the total variation distance that
$||\mathcal{L}_{Z_{1}|Z_{0}=z_{0}}-\mathcal{L}_{X_{1}|X_{0}=z_{0}}||_{TV}\leq||\mathcal{L}_{(\phi(Z_{0}),Z_{1})|Z_{0}=z_{0}}-\mathcal{L}_{(\phi(X_{0}),X_{1})|X_{0}=z_{0}}||_{TV}.$
Additionally, since $\xi_{Z}$ and $\xi_{X}$ both have the same mean $m$ and variance $\sigma^{2}$, finite third moments, and lattice size one, we know from [26, Theorem 9] that, for a given $u\in\mathbbm{N}_{1}$, there exists a constant $c$ depending on $\xi_{Z}$ and $\xi_{X}$ such that
$||\mathcal{L}_{Z_{1}|\phi(Z_{0})=u}-\mathcal{L}_{X_{1}|\phi(Z_{0})=u}||_{TV}=||\mathcal{L}_{\sum_{i=1}^{u}\xi_{Z,i}}-\mathcal{L}_{\sum_{i=1}^{u}\xi_{X,i}}||_{TV}\leq\frac{c}{\sqrt{u}}.$
@@
$\quad||\mathcal{L}_{(\phi(Z_{0}),Z_{1})|Z_{0}=z_{0}}-\mathcal{L}_{(\phi(X_{0}),X_{1})|X_{0}=z_{0}}||_{TV}$
$\leq||\mathcal{L}_{\phi(Z_{0})|Z_{0}=z_{0}}-\mathcal{L}_{\phi(X_{0})|X_{0}=z_{0}}||_{TV}+\mathbbm{P}(\phi_{Z}(Z_{0})\leq N|Z_{0}=z_{0})+\frac{c}{\sqrt{N+1}}$
$=||\mathcal{L}_{\phi(z_{0})}-\mathcal{L}_{\phi(z_{0})}||_{TV}+\mathbbm{P}(\phi(z_{0})\leq N)+\frac{c}{\sqrt{N+1}}$
@@
$||\mathcal{L}_{(\phi(Z_{0}),Z_{1})|Z_{0}=z_{0}}-\mathcal{L}_{(\phi(X_{0}),X_{1})|X_{0}=z_{0}}||_{TV}$ $\leq\frac{\nu^{2}(z_{0})}{\big{(}\varepsilon(z_{0})-\lfloor\alpha\cdot\varepsilon(z_{0})\rfloor\big{)}^{2}}+\frac{c}{\sqrt{\lfloor\alpha\cdot\varepsilon(z_{0})\rfloor+1}}$

### v2

@@ 5 Proofs of non-existence results
$\mathcal{L}_{(Z_{j+1},\dots,Z_{k})|Z_{j}=u_{j}}-\mathcal{L}_{(X_{j+1},\dots,X_{k})|X_{j}=u_{j}}||_{TV}\leq\boldsymbol{\mathfrak{K}}_{k-j}(u_{j}),$
@@
[…]
for $u_{j-1}\geq M_{3}$ and $M_{3}\leq N<\varepsilon(u_{j-1})\cdot m$. Given (6) and that $\mathcal{L}_{Z_{1}|Z_{0}=z}-\mathcal{L}_{X_{1}|X_{0}=z}||_{TV}=O(z^{-q})$, and taking $N:=\lfloor\alpha t\cdot u_{j-1}\rfloor$ for $\alpha\in(1/t,1)$, we can simplify the above bound to
[…]
@@
$\leq\frac{s}{u_{j-1}^{q}}+\frac{(a\sigma^{2}+bm^{2})}{(1-\alpha)^{2}t^{2}\cdot u_{j-1}}+\boldsymbol{\mathfrak{K}}_{k-j}(\alpha t\cdot u_{j-1})\;=:\boldsymbol{\mathfrak{K}}_{k-j+1}(u_{j-1}).$
Since we assumed that $\boldsymbol{\mathfrak{K}}_{k-j}$ was a decreasing function on $M_{3},\infty)$, we see that $\boldsymbol{\mathfrak{K}}_{k-j+1}$ is also a decreasing function on $M_{3},\infty)$.
Having shown the recursive relationship (17), and since $\alpha\in(1/t,1)$ implies $\alpha t>1$, by taking $f_{i}=\boldsymbol{\mathfrak{K}}_{i}$ for $1\leq i\leq k$ we see that the sequence of functions $\{\boldsymbol{\mathfrak{K}}_{i}\}_{1\leq i\leq k}$ satisfies the requirements of Lemma 5.2. It hence follows that $\lim_{z\to\infty}\lim_{k\to\infty}\boldsymbol{\mathfrak{K}}_{k}(z)=0$. Since for any $z_{0}$,
[…]
Proof of Proposition 3.2..
For a given set $\Pi$ of supercritical CBPs satisfying (6) and with transition probabilities parameterised by $\boldsymbol{\theta}$, assume there exist processes $\{Z_{n},z_{0}\},\{X_{n},z_{0}\}\in\Pi$ with $\boldsymbol{\theta}_{Z}\neq\boldsymbol{\theta}_{X}$ such that $\mathcal{L}_{Z_{1}|Z_{0}=z_{0}}-\mathcal{L}_{X_{1}|X_{0}=z_{0}}||_{TV}=O(z_{0}^{-q})$ for some $q>0$. Let $\{\hat{\boldsymbol{\theta}}_{k}\}_{k\in\mathbbm{N}_{1}}$ be a sequence of estimators for $\boldsymbol{\theta}$.
Suppose that the sequence of estimators $\{\hat{\boldsymbol{\theta}}_{k}\}_{k\in\mathbbm{N}_{1}}$ is (weakly) consistent for $\boldsymbol{\theta}$ on the set of unbounded growth of the process. Then there exists a subsequence $\{k_{j}\}_{j\in\mathbbm{N}_{1}}$ such that $\{\hat{\boldsymbol{\theta}}_{k_{j}}\}_{j\in\mathbbm{N}_{1}}$ forms a strongly consistent sequence of estimators on the set of unbounded growth—that is, on that set, $\hat{\boldsymbol{\theta}}:=\lim_{j\to\infty}\hat{\boldsymbol{\theta}}_{k_{j}}$ exists and equals $\boldsymbol{\theta}$ almost surely (see, for example, [6, Theorem 2.3.2]). In a slight abuse of notation, we say that $\hat{\boldsymbol{\theta}}$ is itself strongly consistent.
@@
$\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}=\sup_{C\subseteq\mathbbm{N}_{0}^{\infty}}|\mathbbm{P}(\{Z_{n},z_{0}\}\in C)-\mathbbm{P}(\{X_{n},z_{0}\}\in C)|,$
@@
$\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}\geq|\mathbbm{P}(\hat{\boldsymbol{\theta}}(\{Z_{n},z_{0}\})=\boldsymbol{\theta}_{X},\;Z_{n}\to\infty)-\mathbbm{P}(\hat{\boldsymbol{\theta}}(\{X_{n},z_{0}\})=\boldsymbol{\theta}_{X},\;X_{n}\to\infty)|.$ (18)
But, since $\hat{\boldsymbol{\theta}}$ is a strongly consistent estimator for $\boldsymbol{\theta}$ on the set of unbounded growth, and $\boldsymbol{\theta}_{Z}\neq\boldsymbol{\theta}_{X}$,
$\mathbbm{P}(\hat{\boldsymbol{\theta}}(\{Z_{n},z_{0}\})=\boldsymbol{\theta}_{X},\;Z_{n}\to\infty)=0,\qquad\textrm{while}\qquad\mathbbm{P}(\hat{\boldsymbol{\theta}}(\{X_{n},z_{0}\})=\boldsymbol{\theta}_{X},\;X_{n}\to\infty)=\mathbbm{P}(X_{n}\to\infty).$
@@
$\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}\geq\mathbbm{P}(X_{n}\to\infty).$
@@
[…]
@@
    Given (6), [9, Theorem 3.2] tells us that $\mathbbm{P}(Z_{n}\to\infty)\to 1$ as the initial population size $z_{0}$ approaches infinity.
This creates a contradiction, since (i) and (ii) tell us that we can find $z_{0}\in\mathbbm{N}_{1}$ sufficiently large such that $\mathcal{L}_{\{Z_{n},z_{0}\}}-\mathcal{L}_{\{X_{n},z_{0}\}}||_{TV}<\mathbbm{P}(X_{n}\to\infty)$. So, in conclusion, no consistent estimator for $\boldsymbol{\theta}$ exists on the set of unbounded growth.
@@
Proof of Lemma 3.3..
Since our two processes $\{Z_{n},z_{0}\}$ and $\{X_{n},z_{0}\}$ belong to the set $\Pi^{(k)}$, they must have a common control function, $\phi(\cdot)$. If we consider the realisations of $\phi(\cdot)$ as forming interstitial states within the processes $\{Z_{n},z_{0}\}$ and $\{X_{n},z_{0}\}$, we can recognise the expanded processes $\{\phi(z_{0}),\,Z_{1},\,\phi(Z_{1}),\,Z_{2},\,\dots\}$ and $\{\phi(z_{0}),\,X_{1},\,\phi(X_{1}),\,X_{2},\,\dots\}$, consisting of alternating progenitor numbers and population sizes, as two time-inhomogeneous Markov chains. We can see from the definition of the total variation distance that
[…]
Additionally, since $\xi_{Z}$ and $\xi_{X}$ both have the same mean $m$ and variance $\sigma^{2}$, finite third moments, and lattice size one, we know from [28, Theorem 9] that, for a given $u\in\mathbbm{N}_{1}$, there exists a constant $c$ depending on $\xi_{Z}$ and $\xi_{X}$ such that
[…]
@@
[…]
@@
[…]